# CARE: Confidence-Aware Regression Estimation of building density fine-tuning EO Foundation Models

### v1

@@ Abstract
Performing accurate confidence quantification and assessment is important for deep neural networks to predict their failures, improve their performance and enhance their capabilities in real-world applications, for their practical deployment. For pixel-wise regression tasks, confidence quantification and assessment has not been well addressed in the literature, in contrast to classification tasks like semantic segmentation. The softmax output layer is not used in deep neural networks that solve pixel-wise regression problems. In this paper, to address these problems, we develop, train and evaluate the proposed model Confidence-Aware Regression Estimation (CARE). Our model CARE computes and assigns confidence to regression results. We focus on solving regression problems as downstream tasks of a Foundation Model for Earth Observation (EO). We evaluate the proposed model CARE and experimental results on data from the Copernicus Sentinel-2 satellite constellation for estimating the density of buildings show that the proposed method can be successfully applied to regression problems. We also show that our approach outperforms other methods.
@@
Earth Observation (EO), Foundation Model.
@@ I Introduction
The utility of satellite data seems inexhaustible and, thanks to developments in Artificial Intelligence (AI), applications have begun to emerge at an accelerated pace. The focus is on developing methodological improvements in AI that: i) are needed for further improving real-world practicable solutions that use Earth Observation (EO) data, and ii) are prerequisites of important goals, namely accurately performing tasks that are “downstream” from EO data. The downstream tasks are, for example, building density estimation, land cover classification, crop type mapping and useful insights for Earth actions for sustainability and climate. Learning from unlabeled data, i.e. devising new methods for self-supervised learning, and developing an EO Foundation Model (FM) and evaluation framework with confidence quantification are important.
Self-supervised learning and operating within an EO FM framework are motivated by the lack of labeled data and the technical impossibility to provide labels for the massive volume of data collected by the European Space Agency (ESA) and other institutions. The trove of EO data collected is vast and expanding fast. The Sentinel-2 constellation generates approximately 1.6TB of compressed data daily. The lack of labels stands in the way of accomplishing important downstream tasks that we want to accurately perform. Such tasks that are “downstream” from EO data require the acquisition of labels. Also, because Earth is dynamic, labels change over time.
To use the large amounts of unlabeled data from satellites and achieve label efficiency, i.e. ability to learn from only few labeled data, EO FMs first perform self-supervised learning to extract general representations from unlabeled data and, then, supervised learning to specialize on specific applications, i.e. downstream tasks, by learning from a limited number of labeled data [1, 2]. For FMs, approximately $20\%$ of the labels for the use cases are needed, that would otherwise be required.
For deep neural networks that solve regression problems, confidence estimation is very different from and more challenging than for neural network models that solve classification problems. In addition, many more methods for confidence estimation have been proposed for classification neural networks, rather than for regression neural networks. The problem of accurately assigning a confidence metric to the real-valued output of a regression neural network is challenging and, to the best of the authors’ knowledge, has not been addressed well in the literature. Because the softmax output layer is not available in deep neural networks that solve regression problems, unlike in classification neural networks [4], the interpretation and the definition and meaning of confidence become less clear [6]. To address these problems, we propose a method that computes and assigns a confidence metric to the outputs of regression neural networks. We develop and train the proposed model Confidence-Aware Regression Estimation (CARE). Confidence is a proxy for the probability of correct classification. It is a metric between 0 and 1 that is an indicator of how well do we trust the results of the model. Confidence is an a priori estimate of the performance of our model.
@@
Our main contributions are the development, training and evaluation of the proposed model CARE. We propose a confidence-aware inference method that first sorts the samples in the mini-batch with respect to their errors. In particular, we perform sorting of samples in ascending order, i.e. samples that have low errors go first in the mini-batch, while high errors go last. In this way, the confidence metric takes a value that represents the relative error level [6]. The model is then trained using this proposed mini-batch based training algorithm, where the outputs of the model CARE are both the confidence metric and the regression model output prediction. The desired output for the confidence metric result is based on sorting the samples in the mini-batches based on their error, i.e. absolute error or Mean Squared Error (MSE). Results on multi-spectral Sentinel-2 data, and more specifically for the problem of building density estimation, which is the regression task of predicting how close buildings are to each other, how densely built is the area in the image and how cities expand (i.e. urban growth), have shown that our method is effective.
@@ II Related Work
Many different models exist that perform estimation and confidence assignment. However, most of the existing models are for classification (rather than regression) and use the softmax output probability [6]. Without model retraining, it is possible to correct, as well as calibrate to achieve no overfitting and use the entire range of possible output values, the output softmax probability of deep neural networks [4]. For example, by using temperature scaling, calibration based on the entropy and the difference between the first and second top class probabilities, or correction based on the aggregation of dispersion measures of softmax output probabilities over segments. Moreover, with model retraining, two-headed networks can be used to estimate the True Class Probability (TCP), which is a better estimator of confidence than Maximum Class Probability (MCP) [9].
@@
For regression tasks, estimating model failures and regions of operation where the model should not be trusted is challenging [17]. We note that if we choose to turn a regression task to a classification problem, we lose precision, exactness, the desired outcome of having as an output a real-valued number, and having a result with an infinite number of decimal places. Considering the importance of solving regression problems, for example in EO for canopy height Above Ground Biomass (AGB) estimation or for CO2 flux estimation, it is crucial to develop methods that address the shortcomings of the existing methods and correctly compute and assign a confidence metric to the prediction outputs of regression neural networks [7].
@@
Furthermore, most existing evaluation methods for EO FMs are based only on accuracy. However, EO datasets might have noisy labels and the evaluation of the performance of EO FMs using only the accuracy is not enough and might lead to misleading conclusions. Assessing the confidence with which the model performs inference is essential for usability and uptake. Accuracy, for the models and problems we consider, is not a sufficient performance measure. An evaluation framework for FMs has to include confidence assessment to benchmark and compare different FMs, evaluating their performance on both regression and classification segmentation downstream tasks.
III Our proposed methodology CARE
To increase reliability and trust, accurately computing the confidence with which the model performs inference and outputs its results is desirable for neural networks that solve regression problems. For models to function and be operational in the real-world, estimating and assigning an accurate confidence metric to every output model prediction is important. To successfully improve the performance of the model, assigning a confidence metric to every inference of the model is crucial, as for the correctly estimated low-confidence samples, the model can subsequently improve its performance on them. When we have samples that have low confidence, we need to perform further model retraining on these samples or collect more data near these samples, in this region in the data space. Using the model CARE, we are able to achieve improved performance by assigning a confidence metric to the output results, and this is one of the main reasons that confidence is important. Also, assigning a confidence metric to every output model prediction is essential, because for the low confidence data, the model can choose to abstain from providing a prediction in these specific cases, rather than outputting an incorrect estimation result.
@@ III Our proposed methodology CARE
The confidence metric helps humans to make correct decisions for critical matters, e.g. for policy. Moreover, to effectively combine possibly contradictory results from different models [6], the assigned confidence metrics can be used to better inform the fusion and integration of the different model outputs. Within a Mixture of Experts (MoE) framework, we can effectively combine different models using the confidence metric. The several different models might have outputs that lead to different findings and conclusions. In this way, we are able to associate a confidence measure to the inference of each of the models in the MoE and, then, combine their outputs to the final model decision using the individual confidences.
@@
Using the confidence metric that we have assigned to the model output, we are also able to perform anomaly detection, as low confidence is an indicator of anomalies. This is important as anomaly detection/ Out-of-Distribution (OoD) detection is challenging, especially when the abnormalities are close to the normal/ in-distribution samples in the image data space [15, 16]. Confidence also enables us to perform confidence-based change detection, which is crucial in scenarios when we do not have labels for the change, as well as when the change is rare, occurring a small percentage of the time ($\leq 5\%$).
@@
Loss function of CARE. The function minimized during training with Stochastic Gradient Descent (SGD) is given by
@@
where we denote the neural network parameters of our model by $\boldsymbol{\theta}_{nn}$. Here, the input image is $\textbf{x}_{i}$, the regression output $y_{i}$, the confidence output $c_{i}$, the ground truth regression value $y_{i}^{*}$ and the ground truth confidence $c_{i}^{*}$. The first loss term, $L_{0}(\textbf{x},\boldsymbol{\theta}_{nn},y,y^{*})$, is the distance measure that is given by
@@
where we compute the MSE and use the batch size $N$. The second loss term, $L_{1}(\textbf{x},\boldsymbol{\theta}_{nn},y,c,y^{*},c^{*})$, is for the estimated confidence, where $\forall i,j$: $c_{i}\geq c_{j}\Leftrightarrow d(y_{i},y_{i}^{*})\leq d(y_{j},y_{j}^{*})$. This second loss term is the similarity metric that is given by
$L_{1}(\textbf{x},\boldsymbol{\theta}_{nn},y,y^{*},c,c^{*})=\dfrac{1}{N}\sum_{i=1}^{N}|y_{i}-y_{i}^{*}|\times||c_{i}-c_{i}^{*}||_{2}^{2}\text{.}$ (4)
Architecture. The proposed model CARE estimates the confidence of the outputs of deep neural networks that solve pixel-wise regression problems. It assigns the confidence of $1$ for low error, and 00 for high error. Using the model CARE, we are able to estimate confidence as an unbiased predictor of the error rate. Here, for the implementation of the proposed methodology, we use the EO FM that we have recently trained, PhilEO [1, 3]. Within a FM framework, we are able to tackle and solve jointly a group of problems (downstream tasks) that are of interest to the EO community [5], [2]. For the training of the EO FM with unlabeled data, we have performed self-supervised learning that learns the correlations between: the Sentinel-2 multi-spectral images, masking and reconstruction, geo-location (longitude and latitude) estimation and climate zone classes. The architecture is a modified U-Net model [1]. For the downstream task of building density estimation, we fine-tune the EO FM, regress how close the buildings are in the image and evaluate the performance of the final model.
@@
The variation in the regression is modelled and captured by the confidence metric of CARE. In this paper, we focus on the task of building density regression. We have trained EO FMs using self-supervised learning, followed by supervised learning, making efficient use of labels [5], [1]. A methodology has been devised to take into account the geo-location longitude and latitude satellite information. As a downstream task, we retrain the model to regress building density. To accurately perform confidence estimation, the network has two output heads, i.e. one for the predicted regression and one for the estimated confidence, in (3) and (4). The latter is a metric indicating the extent to which we can trust the output results of the other head, i.e. predicted regression values [6, 7]. For $c_{i}^{*}$, in every mini-batch, after sorting the samples with respect to the error (i.e. MSE), $\eta=80\%$ of the data are assigned to $c_{i}^{*}=1$, while the remaining to $c_{i}^{*}=0$ [6]. This is a ranking problem, where correct predictions are ranked before incorrect ones. For the hyperparameter $\eta$, the threshold $80\%$ is chosen because we perform model retraining and first set $\lambda=0$ in (2) and, then, set $\lambda>0$ [9], [6]. The results of CARE in Sec. IV, because of the first term in (4), are not sensitive to relatively small changes in $\eta$. During inference, for our decision rule, to predict model failure, when the metric $(1-\text{confidence})$ is more than $\zeta=20\%$ of the regression value result, then we have detected incorrect pixels and CARE chooses to abstain from providing an output rather than predicting incorrect values.
IV Evaluation and Results of CARE
The evaluation of the proposed model CARE is based on the observation that the absolute error of the regression output should be equal to the quantity: $(1-\text{confidence})$. This should be true for every pixel and image. We examine if we learn to both: predict building density, and estimate confidence. We compute the average absolute error between: i) the absolute error of the regression output, and ii) the metric $(1-\text{confidence})$. The mean is over the pixels and the images, and the ideal result is zero. In the evaluation results, we are interested in the error between the regression error and the metric $(1-\text{confidence})$ being very small. Therefore, for the evaluation of the proposed model CARE and the estimated confidence, we use the Mean Absolute Error (MAE), and we also calculate the square Root MSE (RMSE). Furthermore, we compute the median absolute error, i.e. using the median instead of the average, where the median is more robust to outliers than the mean. The median error, in our case, is much lower than the average error.

### v2

@@ Abstract
Performing accurate confidence quantification and assessment in pixel-wise regression tasks, which are downstream applications of AI Foundation Models for Earth Observation (EO), is important for deep neural networks to predict their failures, improve their performance and enhance their capabilities in real-world applications, for their practical deployment. For pixel-wise regression tasks, specifically utilizing remote sensing data from satellite imagery in EO Foundation Models, confidence quantification is a critical challenge. The focus of this research is on developing a Foundation Model using EO satellite data that computes and assigns a confidence metric alongside regression outputs to improve the reliability and interpretability of predictions generated by deep neural networks. To this end, we develop, train and evaluate the proposed Confidence-Aware Regression Estimation (CARE) Foundation Model. Our model CARE computes and assigns confidence to regression results as downstream tasks of a Foundation Model for EO data, and performs a confidence-aware self-corrective learning method for the low-confidence regions. We evaluate the model CARE, and experimental results on multi-spectral data from the Copernicus Sentinel-2 constellation to estimate the building density (i.e. monitoring urban growth), show that the proposed method can be successfully applied to important regression problems in EO. We also show that our model CARE outperforms other methods.
@@
Earth Observation (EO), Pixel-wise regression, EO Foundation Models, Remote sensing, Copernicus Sentinel-2
@@ I Introduction
The significance of confidence quantification and assessment in deep learning, specifically in AI Foundation Models in Earth Observation (EO) that use satellite data, for regression applications is critical. The utility of satellite data seems inexhaustible, and thanks to developments in AI, applications emerge at an accelerated pace in EO Foundation Models using remote sensing data. The trove of EO data is vast and expanding fast. The ESA Sentinel-2 constellation generates approximately 1.6TB of compressed multi-spectral data daily. However, the lack of labels stands in the way of accurately performing important downstream tasks, and confidence quantification is critical to achieve better performance.
@@
The focus of this work is on developing methodological improvements in EO Foundation Models that: i) are needed to further improve performance for practicable real-world solutions that use EO data, and ii) are prerequisites to achieve important goals, namely accurately performing tasks that are ‘downstream’ from remote sensing data. The downstream tasks are, for example, building density estimation, land cover classification, crop type mapping and useful insights for Earth actions for sustainability and climate. Learning from unlabeled data, i.e. devising new methods for self-supervised learning, and developing an EO Foundation Model (FM) with confidence quantification and assessment are important. Confidence is a proxy for the probability of correct results. It is a metric between 00 and $1$ that is an indicator of how well do we trust the output results of the model. Confidence is an a priori estimate of the performance of our model.
The problem of accurately assigning a confidence metric to the real-valued output of a pixel-wise regression neural network in EO Foundation Models to estimate building density from satellite data is challenging and, to the best knowledge of the authors, has not been well addressed in the literature.
Our main contributions. To address these problems, we propose the Confidence-Aware Regression Estimation (CARE) Foundation Model for remote sensing and EO. The model CARE computes and assigns confidence metrics alongside regression outputs in order to improve the reliability and interpretability of predictions generated by deep neural networks. The CARE model is implemented based on a modified U-Net architecture suitable for processing multi-spectral satellite images, which allows it to handle the complex spatial features present in the data. Assigning a confidence metric to every inference of the model is crucial, as for the correctly estimated low-confidence samples in building density estimation from satellite data, our model subsequently implements a self-corrective method and improves its performance on them. CARE uses two heads that compute and assign a regression density measure and a confidence metric to the output of regression neural networks. Using our proposed loss function that comprises a Mean Squared Error (MSE) distance loss and a confidence loss, our model successfully performs self-corrective confidence-aware regression for building density estimation which is a downstream task of EO Foundation Models, and outperforms other methods.
@@
Our main contributions are the development, training and evaluation of the proposed CARE model. Results on multi-spectral Sentinel-2 data, and more specifically for the problem of building density estimation, which is the regression task of predicting how close buildings are to each other, how densely built is the area in the image and how cities expand (i.e. urban growth) using EO Foundation Models, have shown that our method is effective and outperforms other base models.
@@ II Related Work
Foundation Models in EO. To use the vast amounts of unlabeled satellite data and achieve label efficiency, i.e. the ability to learn from only few labeled data in fine-tuning, EO Foundation Models are first pretrained on large unlabeled satellite datasets by performing self-supervised learning to extract special representations from unlabeled data [2, 33]. Then, in the downstream applications, Foundation Models are fine-tuned by performing supervised learning to specialize in specific applications, i.e. downstream tasks, by learning from a limited number of labeled data. In inference, classification or regression is performed for the downstream task [1, 2].
Self-supervised learning and operating within an EO FM framework [24, 23] are motivated by the lack of labeled data and the technical impossibility to provide labels for the massive volume of data collected by the European Space Agency (ESA) and other space agencies (e.g., NASA, JAXA). Enhancing EO Foundation Models with confidence quantification and assessment is critical to improve their performance.
Confidence-aware models. There are many different models that perform estimation and confidence assignment [7, 26, 28]. However, most of the existing models are not for pixel-wise regression building density estimation. Existing models for classification are based on the softmax output probability [6]. For regression problems where the result is a real-valued number, the output softmax layer is not used in the neural network [21]. Therefore, for regression neural networks that do not use softmax, it is not straightforward to apply softmax-probability-based classification confidence methods.
@@
Without model retraining, it is possible to correct, as well as calibrate to achieve no overfitting and use the entire range of possible output values, the output softmax probability of deep neural networks [4]. For example, by using temperature scaling: Calibration based on entropy and the difference between the first and second top class probabilities, or correction based on the aggregation of dispersion measures of softmax output probabilities over segments. Moreover, with model retraining, two-headed networks can be used to estimate the True Class Probability (TCP), which is a better confidence estimator than Maximum Class Probability (MCP) [9].
@@
For regression tasks, estimating model failures and regions of operation where the model should not be trusted is challenging [17]. We note that if we choose to turn a regression task into a classification problem [26], then we lose precision, exactness and the desired outcome of having as an output a real-valued number.
Considering the importance of solving regression problems [21], for example in EO for the accurate estimation of canopy height Above Ground Biomass (AGB) [2], or for the estimation of CO2 flux, it is crucial to develop methods that address the shortcomings of existing methods and correctly compute and assign a confidence metric to the prediction outputs of regression neural networks [7].
The work in [19] proposes a spatial self-corrective learning framework which: a) is able to locate predictions with potential large errors by estimating confidence using Maximum Likelihood Estimation (MLE), b) uses confidence-based pseudo-interpolation to correct low-confidence predictions in local neighborhoods, and c) performs recurrent self-refinement of the estimated height in an iterative manner. Experiments on different landscapes in the high-latitude regions show improvements compared to the other baseline methods.
The work in [20] presents a probabilistic regression method. The deep neural network estimates the unnormalized density from the input-target pair using a conditional target density $p(y|x)$ model, with energy as the basis $(x,y)$. Monte Carlo sampling is used to reduce the negative log probability of this probabilistic $p(y|x)$ model.
Furthermore, in the literature [20, 29, 30, 31], in addition to probabilistic models and Bayesian neural networks, for confidence assignment and assessment in regression neural networks, evidential regression can also be performed [35, 36, 37], as well as conformal prediction (i.e. conformal inference) for regression [38, 39].
@@
EO data noisy labels. Most existing evaluation methods for EO Foundation Models [2, 40] are based only on accuracy. However, remote sensing datasets might have noisy labels, i.e. incorrect annotations, and the evaluation of the performance of EO FMs using only the accuracy is not enough and might lead to misleading conclusions. Assessing the confidence with which the model performs inference is essential for usability and uptake [32, 25]. Accuracy, on its own, for the models and problems we consider, is not a sufficient performance measure. An evaluation framework for FMs has to include confidence assessment to benchmark and compare different Foundation Models, evaluating their performance on both regression and classification segmentation downstream tasks.
@@
Confidence metric. The confidence metric helps humans to make correct decisions for critical matters [9, 27, 22], e.g. for policy. Moreover, to effectively combine possibly contradictory results from different models [6], the assigned confidence metrics can be used to better inform the integration and fusion of the different model outputs. Within a Mixture of Experts (MoE) framework, we can effectively combine different models using the confidence metric. Several different models might have outputs that lead to different findings and conclusions. In this way, we are able to associate a confidence measure to the inference of each of the models in the MoE and, then, combine their outputs to the final model decision using the individual confidences.
@@
Using the confidence metric that we have assigned to the model output, we are also able to perform anomaly detection, as low confidence is an indicator of anomalies. This is important as anomaly detection/ Out-of-Distribution (OoD) detection is challenging, especially when the abnormalities are close to the normal/ in-distribution samples in the image data space [15, 16]. In addition, the confidence metric also enables us to perform confidence-based change detection, which is crucial in scenarios when we do not have labels for the change, as well as when the change is rare, occurring only a small percentage of the time (e.g., $<5\%$, or even $<2\%$).
Figure 1: CARE for confidence-aware regression: assign confidence metric to predictions, identify wrong predictions and refine the model.
III Our proposed methodology
Confidence-aware pixel-wise regression. To increase reliability and trust, accurately computing the confidence with which the model performs inference and outputs its results is desirable for neural networks that solve regression problems. For models to function and be operational in the real-world, estimating and assigning an accurate confidence metric to every output model prediction is important. To successfully improve the performance of the model, assigning a confidence metric to every inference output of the model is crucial, as for the correctly estimated low-confidence samples, the model can subsequently improve its performance specifically on them. When we have samples with low confidence, we need to perform further retraining of the model on these samples or collect more data near these samples (or data augmentation), in this region in the data space. Using the model CARE, we are able to achieve improved performance by assigning a confidence metric to the output results, and this is one of the main reasons that confidence is important. Also, assigning a confidence metric to every output model prediction is essential, because for the low confidence data, the model can choose to abstain from providing a prediction in these specific cases, rather than outputting an incorrect estimation result.
III-A Proposed method
The core methodology of the model CARE is based on a mini-batch training algorithm that takes advantage of the computed MSE distance measure assigned to the regression predictions. CARE uses two heads that compute and assign a regression building density measure and a confidence metric to the output of regression neural networks. The approach is built upon sorting samples within mini-batches based on their prediction absolute errors, which allows for establishing a relative confidence metric for each sample.
Sorting mechanism. In each mini-batch, samples are sorted in an ascending order of their errors (using the MSE as a measure of error). This enables the model to assign higher confidence scores to samples with lower errors and vice versa.
Training procedure with dual outputs. The model operates with two output heads: one for the regression prediction and another for the confidence metric. The training of the model aims to minimize a custom loss function that consists of two components: a) Regression Loss (i.e. $L_{0}$): Measures the distance between the predicted outputs and the ground truth values using the MSE. b) Confidence Loss ($L_{1}$): Assesses the relationship between the prediction absolute errors and the assigned confidence scores, ensuring that higher confidence correlates with lower prediction errors.
@@ III-A Proposed method
Loss function. The function minimized during training with Stochastic Gradient Descent (SGD) is given by:
@@
where we denote the neural network parameters of our model by $\boldsymbol{\theta}_{nn}$. In the combined loss function (1), $\lambda$ is a hyperparameter that can be adjusted to weigh the importance of confidence estimation relative to regression accuracy during training. Here, the input image is $\textbf{x}_{i}$, the regression output $y_{i}$, the confidence output $c_{i}$, the ground truth regression value $y_{i}^{*}$ and the ground truth confidence $c_{i}^{*}$. The first loss term, $L_{0}(\textbf{x},\boldsymbol{\theta}_{nn},y,y^{*})$, is the distance measure that is given by:
@@
where we compute the MSE and use the batch size $N$. The second loss term, $L_{1}(\textbf{x},\boldsymbol{\theta}_{nn},y,c,y^{*},c^{*})$, is for the estimated confidence, where $\forall i,j$: $c_{i}\geq c_{j}\Leftrightarrow d(y_{i},y_{i}^{*})\leq d(y_{j},y_{j}^{*})$. This second loss term is the confidence metric that is given by:
$L_{1}(\textbf{x},\boldsymbol{\theta}_{nn},y,y^{*},c,c^{*})=\dfrac{1}{N}\sum_{i=1}^{N}|y_{i}-y_{i}^{*}|\cdot||c_{i}-c_{i}^{*}||_{2}^{2}\text{.}$ (4)
Architecture. The model CARE is implemented based on a modified U-Net architecture that has approximately $70$M parameters, suitable for processing multi-spectral satellite images in EO Foundation Models, which allows it to handle the complex spatial features present in the data [1]. The U-Net’s encoder-decoder structure is particularly adept at tasks requiring precise localization, important for pixel-wise predictions in EO tasks.
Datasets. Our model CARE is pretrained on the PhilEO Globe Sentinel-2 global dataset [1, 3] and fine-tuned on the PhilEO Bench downstream tasks dataset [1]. CARE and the other examined baseline models in this work are evaluated on the PhilEO Bench [1, 3, 4, 5].
III-B CARE flowchart diagram
The proposed model CARE estimates the regression value and the confidence metric in the output of the deep neural network that solves pixel-wise regression problems in Fig. 1. CARE assigns the confidence of $1$ for low error, and 00 for high error.
For the training of the EO Foundation Model with unlabeled data, we have performed self-supervised learning that learns the correlations between: the Sentinel-2 multi-spectral images, masking and reconstruction, geolocation (longitude and latitude) estimation and climate zone classes. For the downstream task of building density estimation, we fine-tune the EO Foundation Model, regress how close the buildings are in the image and evaluate the performance of the final model.
In this paper, we focus on the task of building density regression. To accurately perform confidence estimation, the network has two output heads, i.e. one for the predicted regression output value and one for the estimated confidence, in (3) and (4). The latter is a confidence metric that correlates with the ground truth, indicating the extent to which we can trust the output results of the other head, i.e. the predicted regression values [6, 7].
For $c_{i}^{*}$, in every mini-batch, after sorting the samples in an ascending order with respect to their errors (i.e. MSE), $\eta=80\%$ of the data are assigned with high confidence scores (i.e. $c_{i}^{*}=1$), while the remaining data to low confidence scores (i.e. $c_{i}^{*}=0$). For the hyperparameter $\eta$, the threshold $80\%$ is chosen because we perform model retraining, and we first set $\lambda=0$ in (2) and, then, set $\lambda>0$ [9], [6]. During inference, for our decision rule, to predict model failure, when the metric $(1-\text{confidence})$ is more than $\zeta=20\%$ (or $10\%$) of the regression value result, then we have detected incorrect pixels and CARE performs self-correction by choosing to abstain from providing an output rather than predicting incorrect regression values.
@@ III-B CARE flowchart diagram
IV Evaluation and Results
The evaluation of the proposed model CARE is based on the observation that the absolute error of the regression output should be equal to the quantity: $(1-\text{confidence})$. This should be true for every pixel and image. We examine if we learn to both: predict building density, and estimate confidence. We compute the average absolute error between: i) the absolute error of the regression output, and ii) the metric $(1-\text{confidence})$. The mean is over the pixels and the images, and the ideal result is zero. In the evaluation results, we are interested in the error between the regression absolute error and the metric $(1-\text{confidence})$ being very small. Therefore, in this work, for the evaluation of the proposed model CARE and the estimated confidence, we use the MSE. Furthermore, we compute the median absolute error, i.e. using the median instead of the average, where the median is more robust to outliers than the mean. The median absolute error, in our case for building density estimation, is much lower than the average error.
IV-A Evaluation of CARE using absolute error in Table I and comparison to other models
In Table I, we evaluate the proposed model CARE and examine both the average error and the median error for the results of the model. The median error of the proposed model CARE is low, i.e. $0.00184$, and this is desirable as it shows that the difference between the absolute error of the regression output and the estimated metric $(1-\text{confidence})$ is small. We also evaluate the regression output result using the MSE, in Tables I and II. We run experiments using $n$-sample tests, where we use $n$ training samples per region for fine-tuning. Stratified sampling is performed, and the different regions are: Denmark, East Africa, Egypt, Guinea, Europe, Ghana, Israel, Japan, Nigeria, North America, Senegal, South America, Tanzania and Uganda [1, 5]. The performance of the model CARE, for $n=5000$, is examined in Table I.